Thank you God!!!
Ft=mv-mu
F=42
t=5
U=10
M=7
Find “v”
Can anyone please answer my question??? With the simplest working, like standard 6 workings , thanks a lot

Answer by Alan3354(69443)

Ft=mv-mu
F=42
t=5
U=10
M=7
Find “v”
=============
Ft + mu = mv
v = (Ft - mu)/m
v = (42*5 - 7*10)/7
v = (210 - 70)/7
v = 20